# The Riemann Hypotheses is false

### v22

@@
say. For $0<a<b\leq 2a$ and $t\in\mathbb{R}_{>0}$, note that
@@
Appendix 1. Proof of (10) from (6):
@@
For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}.$ (22)
By partial summation, it follows from (6) that
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|(\log T)^{k+1}}.$ (23)
Inserting (23) into the right-hand side of (22) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$ gives
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (24)
as desired, for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎

### v24

@@
say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
@@
Appendix 1. Derivation of (10) from (6):
@@
For $T>n\geq m+1\geq\frac{T}{\log^{2}T}$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=m^{-it}\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}.$ (22)
Define $c_{n}:=n^{it},C(x):=\sum_{n\leq x}c_{n}$ and $f(n)=\frac{1}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$. By partial summation, note that
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ $=\sum_{m+1\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ (23)
$=\int_{m+1}^{T}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (24)
$=\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (25)
$\ll T^{1/2}\log^{4}T\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}|f^{\prime}(x)|\mathrm{d}x$ (26)
$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log^{k+1}(\frac{T}{\log^{2}T})}$ (27)
where (26) was obtained from (25) by applying (6), which entails that $C(x)-C(m+1)\ll T^{1/2}\log^{4}T$ uniformly for $m+1\leq x\leq T$. Inserting (26) into the right-hand side of (22) and invoking the fact that
$\frac{1}{\log(T/\log^{2}T)}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log(T/\log^{2}T)}\Big{)}^{k}=\frac{1}{\log(T/m\log^{2}T)},$ (28)
gives
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(\frac{T}{m\log^{2}T})}$ (29)
$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (30)
as desired. This completes the proof for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎